OpenStudy (anonymous):

The sum of a number and 9 is -2. a. n - 9 = -2; 7 c. -n + 9 = -2; 11 b. n + 9 = -2; -11 d. n + 9 = 2; -7

OpenStudy (usukidoll):

sum means addition.

OpenStudy (wolf1728):

So, continuing with UsukiDoll's answer we can rephrase your question to sum of a number + 9 = -2 need more help?

OpenStudy (anonymous):

n+9= -2 n +9-9 = -2-9 therefore n = -11

OpenStudy (anonymous):

this is not direct answer !!

TheSmartOne (thesmartone):

Hi @yajna , On OpenStudy we have a set of rules called the Code of Conduct that each user must follow to avoid getting in trouble with the moderators. When you signed up, you agreed to those rules. Please make sure to read the Code of Conduct: http://openstudy.com/code-of-conduct One of the rules is to guide the asker and not give them the answer. Make sure to guide them as OpenStudy values the learning process. Thanks :)

TheSmartOne (thesmartone):

Yajna, your reply didn't engage the asker in any way. By giving them the steps and the answer it does not help them learn. OpenStudy was made so that a person could learn from their peers how to solve a question that was difficult for them but easy for someone else. So keep in mind from next time to guide the asker. Give them hints on how to solve their question -- don't do it for them.